OpenStudy (anonymous):

Divide. –7.2 ÷ –0.9 A. –8 B. –0.8 C. 0.8 D. 8

OpenStudy (tkhunny):

What's preventing you from doing it? \(\dfrac{-7.2}{-0.9} = \dfrac{7.2}{0.9} = \dfrac{72}{9}\) It just keeps getting more and more fun! The art of writing equivalent fractions will serve you well. Get good at it!